# Dual Quaternion SE(3) Synchronization with Recovery Guarantees

### v1

@@ Appendix C Proofs in Section 2
$\operatorname{d}(\bm{w}^{k},\bm{v}_{1})$ $=\min_{z\in\mathrm{UDQ}}\|\bm{w}^{k}-\bm{v}_{1}z\|_{2}\leq\left\|\bm{w}^{k}-\bm{v}_{1}\alpha_{1}{|\alpha_{1}|}^{-1}\right\|_{2}$ (D.10)
$\leq\|\bm{v}_{1}\|_{2}O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}$
$\qquad+\sum_{j=2}^{n}\|\bm{v}_{j}\|_{2}|\lambda_{j}|^{k}|\alpha_{j}|\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}$
$=\Big[\sum_{j=2}^{n}|\lambda_{j}|^{k}|\alpha_{j}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,st}|^{2}/|\lambda_{1,st}|)^{k}\bigr)\Big]^{-1}.$
We use $\sqrt{n}\bm{w}^{k}$ as an initializer for Algorithm 2. As in the proof of Theorem 3.2, it suffices to enforce
$\operatorname{d}_{st}(\bm{w}^{k},\bm{v}_{1})\leq\frac{1}{35},\qquad\operatorname{d}_{\mathcal{I}}(\bm{w}^{k},\bm{v}_{1})\leq\frac{1}{25},$ (D.11)
to guarantees that $\sqrt{n}\bm{w}^{k}$ enters the basin of attraction with
$\operatorname{d}_{st}(\sqrt{n}\bm{w}^{k},\hat{\bm{x}})\leq\operatorname{d}_{st}(\sqrt{n}\bm{w}^{k},\sqrt{n}\bm{v}_{1})+\operatorname{d}_{st}(\sqrt{n}\bm{v}_{1},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}$
and
$\operatorname{d}_{\mathcal{I}}(\sqrt{n}\bm{w}^{k},\hat{\bm{x}})\leq\operatorname{d}_{\mathcal{I}}(\sqrt{n}\bm{w}^{k},\sqrt{n}\bm{v}_{1})+\operatorname{d}_{\mathcal{I}}(\sqrt{n}\bm{v}_{1},\hat{\bm{x}})\leq\frac{\sqrt{n}}{18}.$
By the dual quaternion calculation rules, for all sufficiently large $k$ there exist real constants $M_{{}_{st}},M_{\mathcal{I}},N_{1},N_{2}$, such that
$\operatorname{d}_{st}(\bm{w}^{k},\bm{v}_{1})=\frac{|\lambda_{2,st}|^{k}|\alpha_{2,st}|+M_{st}|\lambda_{2,st}|^{k}}{|\lambda_{1,st}|^{k}|\alpha_{1,st}|+M_{st}|\lambda_{2,st}|^{k}},$ (D.12)
and
$\operatorname{d}_{\mathcal{I}}(\bm{w}^{k},\bm{v}_{1})=\frac{N_{2}|\lambda_{2,st}|^{k}+M_{\mathcal{I}}|\lambda_{2,st}|^{k}}{N_{1}|\lambda_{1,st}|^{k}+M_{\mathcal{I}}|\lambda_{2,st}|^{k}}.$ (D.13)
Specifically,
$M_{st}=M_{st}(\alpha_{1,{st}},\lambda_{1,{st}},\lambda_{1,\mathcal{I}}),\ \ \ M_{\mathcal{I}}=M_{\mathcal{I}}(\alpha_{1,{st}},\alpha_{1,\mathcal{I}},\lambda_{1,{st}},\lambda_{1,\mathcal{I}}),$
$N_{1}=N_{1}(\lambda_{1,{st}},\lambda_{1,\mathcal{I}},\alpha_{1,{st}}),\ \ \ \ \ \ N_{2}=N_{2}(\lambda_{2,{st}},\lambda_{2,\mathcal{I}},\alpha_{2,{st}}).$
Let $r:=|\lambda_{1,_{st}}/\lambda_{2,_{st}}|>1$. Solving (D.12)–(D.13) under (D.11) yields
$K_{\text{init}}\geq\max\left\{\log_{r}\left({\frac{35|\alpha_{2,st}|+34M_{st}}{|\alpha_{1,st}|}}\right),\log_{r}\left({\frac{25N_{2}+24M_{\mathcal{I}}}{N_{1}}}\right)\right\}.$
@@
$\operatorname{d}_{st}(\bm{x}^{0},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25},\quad\operatorname{d}_{\mathcal{I}}(\bm{x}^{0},\hat{\bm{x}})\leq\frac{\sqrt{n}}{18}.$
Substituting these bounds into the estimation error recursive formulas derived in Theorem 3.2 (specifically Eq. (3.1) and Eq. (3.2)) directly yields the claimed error bounds for $\bm{x}^{k}$. This completes the proof.

### v2

@@ Appendix C Proofs in Section 2
$\operatorname{d}(\bm{w}^{k},\bm{v}_{1})$ $=\min_{z\in\mathrm{UDQ}}\|\bm{w}^{k}-\bm{v}_{1}z\|_{2}\leq\left\|\bm{w}^{k}-\bm{v}_{1}\alpha_{1}{|\alpha_{1}|}^{-1}\right\|_{2}$ (35)
$\leq\|\bm{v}_{1}\|_{2}O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}$
$\quad+\sum_{j=2}^{n}\|\bm{v}_{j}\|_{2}|\lambda_{j}|^{k}|\alpha_{j}|\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}$
$=\Big[\sum_{j=2}^{n}|\lambda_{j}|^{k}|\alpha_{j}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]\Big[|\lambda_{1}|^{k}|\alpha_{1}|+O_{D}\bigl((|\lambda_{2,\mathrm{st}}|^{2}/|\lambda_{1,\mathrm{st}}|)^{k}\bigr)\Big]^{-1}.$
We use $\Pi(\sqrt{n}\bm{w}^{k})$ as an initializer for Algorithm 2. As in the proof of Theorem 3.2, it suffices to enforce
$\operatorname{d}_{\mathrm{st}}(\bm{w}^{k},\bm{v}_{1})\leq\frac{1}{70}$ (36)
to guarantee that $\Pi(\sqrt{n}\bm{w}^{k})$ enters the basin of attraction with
$\operatorname{d}_{\mathrm{st}}(\Pi(\sqrt{n}\bm{w}^{k}),\hat{\bm{x}})\leq 2\operatorname{d}_{\mathrm{st}}(\sqrt{n}\bm{w}^{k},\sqrt{n}\bm{v}_{1})+\operatorname{d}_{\mathrm{st}}(\sqrt{n}\bm{v}_{1},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}.$
By the dual quaternion calculation rules, for all sufficiently large $k$, there exists $M_{{}_{\mathrm{st}}}>0$ depending on $\{\alpha_{j},\lambda_{j}\}_{j=1}^{n}$, such that for all sufficiently large $k$
$\operatorname{d}_{\mathrm{st}}(\bm{w}^{k},\bm{v}_{1})\leq\frac{|\lambda_{2,\mathrm{st}}|^{k}|\alpha_{2,\mathrm{st}}|+M_{\mathrm{st}}|\lambda_{2,\mathrm{st}}|^{k}}{|\lambda_{1,\mathrm{st}}|^{k}|\alpha_{1,\mathrm{st}}|+M_{\mathrm{st}}|\lambda_{2,\mathrm{st}}|^{k}}.$ (37)
Let $r:=|\lambda_{1,_{\mathrm{st}}}/\lambda_{2,_{\mathrm{st}}}|>1$. Solving (37) under (36) yields
$K_{\text{init}}\geq\log_{r}\left({\frac{70|\alpha_{2,\mathrm{st}}|+69M_{\mathrm{st}}}{|\alpha_{1,\mathrm{st}}|}}\right).$
@@
$\operatorname{d}_{\mathrm{st}}(\bm{x}^{0},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}.$
Substituting these bounds into the estimation error recursive formulas derived in Theorem 3.2 (specifically Eq. (6) directly yields the claimed error bounds for $\bm{x}^{k}$. This completes the proof.
Appendix E Remarks on Assumptions in Theorem 3.2
Condition (7) is a bounded aligned translation assumption. Indeed, under the unit-dual-quaternion representation
$r_{i}^{k}=\left(1+\frac{\epsilon}{2}\hat{t}_{i}^{k}\right)\hat{q}_{i}^{k},$
we have
$|(r_{i}^{k})_{\mathcal{I}}|=\frac{1}{2}\|\hat{t}_{i}^{k}\|_{2}.$
Then (7) is equivalent to
$\max_{k,i}\|\hat{t}_{i}^{k}\|_{2}\leq 2B.$
Thus, $B$ controls the radius of the ground truth translations after the global alignment $z_{k}$.
Condition (8) is a local nondegeneracy condition for the normalization map. The dual part of the normalization $\mathcal{N}(\bm{y})$ contains a factor $|\bm{y}_{\mathrm{st}}|^{-1}$. Hence the normalization is uniformly Lipschitz in its dual component only when $|\bm{y}_{\mathrm{st}}|$ is bounded away from zero. In the DQGPM step, the normalization input is $\frac{1}{n}\bm{C}\bm{x}^{k}$. In the noiseless case and at the aligned ground truth, namely when $\bm{C}=\hat{\bm{x}}\hat{\bm{x}}^{*}$ and $\bm{x}^{k}=\hat{\bm{x}}z_{k}$, we have
$\frac{1}{n}\bm{C}\bm{x}^{k}=\frac{1}{n}\hat{\bm{x}}\hat{\bm{x}}^{*}\hat{\bm{x}}z_{k}=\hat{\bm{x}}z_{k}=\bm{x}^{k},$
whose standard component has unit norm entrywise. Therefore the threshold $\frac{1}{2}$ simply requires the iterates to remain in a neighborhood where the normalization map is nonsingular.
Condition (9) controls the effect of changing the lexicographic aligner from $z_{k}$ to $z_{k+1}$. Since the aligner is chosen according to the lexicographic dual-number distance, a change of alignment may transfer standard mismatch into the dual component. The constant $\gamma$ quantifies this leakage. It does not need to be small for the contraction argument; it only appears in the additive forcing terms of the dual-component recursion.
Appendix F Additional Experiments on Heavy-tailed Perturbations and Scalability
We further test robustness under heavy-tailed perturbations by replacing the Gaussian model above with a Gaussian-mixture noise model on the observed edges. Specifically, with probability $1-\eta$, an observed relative motion is corrupted by the base perturbation level $(\sigma_{t},\sigma_{r})=(0.1,10^{\circ})$, while with probability $\eta$, it is corrupted by a larger perturbation $(\sigma_{t},\sigma_{r})=(0.4,40^{\circ})$. $\mathrm{error}_{r}$ and $\mathrm{error}_{t}$ are reported in Figure 6 with varying outlier proportion $\eta$ under $p\in\{0.05,0.3\}$. We also fix $\eta=0.2$ and vary the outlier perturbation level by changing $\sigma_{r}$, with $\sigma_{t}=0.01\sigma_{r}$ for each level, as shown in Figure 7. In all settings, DQGPM remains the most accurate method among DQGPM, SPEC, and EIG. These results indicate that DQGPM is robust to heavy-tailed perturbations; handling completely arbitrary gross outliers would require an additional robust reweighting or trimming layer. Finally, to examine scalability, we fix $p=0.1$ and the base noise level $(\sigma_{t},\sigma_{r})=(0.1,10^{\circ})$, and vary the problem size $n$. The CPU times in Figure 8 show that DQGPM scales favorably and remains substantially faster than SPEC, while being competitive with EIG.
(a) Rotation error ($p=0.05$)
(b) Translation error ($p=0.05$)
(c) Rotation error ($p=0.30$)
(d) Translation error ($p=0.30$)
Figure 6: Rotation and translation errors vs. outlier proportion under Gaussian-mixture heavy-tail noise, with base perturbation level $(\sigma_{t},\sigma_{r})=(0.1,10^{\circ})$ and outlier perturbation level $(\sigma_{t},\sigma_{r})=(0.4,40^{\circ})$.
(a) Rotation error ($p=0.05$)
(b) Translation error ($p=0.05$)
(c) Rotation error ($p=0.30$)
(d) Translation error ($p=0.30$)
Figure 7: Rotation and translation errors vs. outlier $\sigma_{r}$ under Gaussian-mixture heavy-tail noise, with $\sigma_{t}=0.01\,\sigma_{r}$ for each outlier perturbation, base perturbation level $(\sigma_{t},\sigma_{r})=(0.1,10^{\circ})$, and outlier proportion $0.2$.
Figure 8: CPU time versus problem scale $n$ under observation rate $p=0.1$ and base noise level $(\sigma_{t},\sigma_{r})=(0.1,10^{\circ})$.